Scenario A thin hoop of mass M and radius R is released from rest at the top of a ramp of length L as shown at right. The ramPARTC: Derive an expression for the speed of the center of mass of the hoop when it reaches the bottom of the ramp.
Answer #1

h = L sino 2 Using conservation of the energy 1 mgh = mv + Int Mere I = MR² , w = So mg h = m + 1 - mig h = mv + mv samy = mg